If you have a weak acid such as acetic acid (pka=4.0) at which pH ranges would you expect it to be a good buffer?
Answer #1

Every buffer has some buffer capacity.

The general range of buffer is pKa±1

i.e lower range is pka -1 and upper range is pka+1

given pKa=4

lower range =4-1 =3 ; upper range pka+1 =4+1 = 5

so it act as buffer in the ph range of 3.0 - 5.0